OpenStudy (anonymous):

Solve the equation 2(h – 8) – h = h – 16 A)h = 8 B)h = –8 C)h = infinitely many solutions D)h = no solution

OpenStudy (radar):

2h-16-h = h -16 0h=0 C will work 0 times anything is still 0.